# Automated threshold selection and associated inference uncertainty for univariate extremes

### v3

@@ 1 Introduction
An inherent challenge in risk modelling is the estimation of high quantiles, known as return levels, beyond observed levels. Such inference is important for designing policies or protections against future extreme events, e.g., in finance or hydrology (Smith,, 2003; Coles et al.,, 2003). Extreme value methods achieve this extrapolation by using asymptotically exact models to approximate the tail of a distribution above a high, within-sample, threshold. They require the selection of a threshold $u$, such that above $u$, extreme value methods provide a suitable model. Here, we develop novel methods for automatic selection of the threshold and propogation of the uncertainty in this selection in return level inferences.
Throughout, we assume that all data are realisations of an independent and identically-distributed (iid) univariate continuous random variable $X$ with unknown distribution function $F$, with upper endpoint $x^{F}:=\text{sup}\{x:F(x)<1\}$. Under weak conditions, Pickands, (1975) shows that for $X>u$, with $u<x^{F}$, the distribution of the rescaled excess $Y=X-u$, converges in distribution to the generalised Pareto distribution (GPD) as $u\rightarrow x^{F}$. To use this limit result in practice, a within-sample threshold $u$ is chosen, above which this limit result is treated as exact. Specifically, whatever the form of $F$, the excesses $Y$ of $u$ are modelled by the single flexible GPD$(\sigma_{u},\xi)$ family, with distribution function
@@
with $y>0$, $w_{+}=\text{max}(w,0)$, $(\sigma_{u},\xi)\in\mathbb{R}_{+}\times\mathbb{R}$ being scale and shape parameters. The exponential distribution arises when $\xi=0$, i.e., as $\xi\rightarrow 0$ in distribution (1), whereas for $\xi>0$, the distribution tail decay is polynomial. For $\xi<0$, $X$ has a finite upper end-point at $u-\sigma_{u}/\xi$ but is unbounded above for $\xi\geq 0$. To estimate the $(1-p)$th quantile, $x_{p}$, of $X$, for $p<\lambda_{u}:=\mathbb{P}(X>u)$, then we can we solve $\hat{F}(x_{p})=1-p$, where $\hat{F}(x_{p})=1-\hat{\lambda}_{u}[1-H(x_{p}-u;\hat{\sigma}_{u},\hat{\xi})]$, where $\hat{\lambda}_{u}$ is the proportion of the realisations of $X$ exceeding $u$, and $(\hat{\sigma}_{u},\hat{\xi})$ are maximum likelihood estimates (MLEs) obtained by using realisations of the threshold excesses. Davison and Smith, (1990) overview the properties of the GPD.
@@
Threshold selection involves a bias-variance trade-off: too low a threshold is likely to violate the asymptotic basis of the GPD, leading to bias, whilst too high a threshold results in very few threshold excesses with which to fit the model, leading to high parameter and return level uncertainty. Thus, we must choose as low a threshold as possible subject to the GPD providing a reasonable fit to the data. There are a wide variety of methods aiming to tackle this problem (Scarrott and MacDonald,, 2012; Belzile et al.,, 2023) with the most commonly used methods suffering from subjectivity and sensitivity to tuning parameters.
@@
A novel, and conceptually much simpler, approach to automated threshold selection than these existing methods, is proposed by Varty et al., (2021) for modelling large, human-induced earthquakes, where data are complex due to improvements in measurement equipment over time. The major implication of such change is that data are missing-not-at-random, with the dataset appearing to be realisations of a non-identically distributed variable, requiring a threshold $u(t)$ which varies with time $t$, even though the underlying process is believed to be identically distributed over $t$. As excesses of $u(t)$ do not have the same GPD parameters over $t$, Varty et al., (2021) transform the excesses to a common standard exponential distribution, via the probability integral transform, using estimates of $(u(t),\sigma_{u(t)},\xi)$ and then, quantify the model fit using a metric based on a QQ-plot and optimise this metric over the time-varying threshold choices. The key novel aspect of their assessment is the use of bootstrapping methods in the metric evaluation which fully accounted for the uncertainty in the GPD fit.
@@
Due to the lack of existing threshold selection methods designed for the context of Varty et al., (2021), the paper focuses on the data analysis rather than investigating the performance of the threshold selection method. We explore how their ideas can be best adapted to threshold selection in an iid context. We find that a variant of the Varty et al., (2021) metric improves the performance and leads to substantially better results than existing automated methods, including greater stability with respect to tuning parameters.
@@
We differ from Varty et al., (2021) as we study both threshold selection and return level estimation when the truth is known. We also address an entirely different problem of how to incorporate the uncertainty resulting from threshold selection into return level estimation. Existing methods typically treat the threshold, once it has been selected, as known, for subsequent return level inference. The available data above candidate threshold choices are often few and so inference can be highly sensitive to the chosen threshold. Reliance on a single threshold leads to poor calibration of estimation uncertainty and as a result, can mislead inference. In particular, we show that the resulting confidence intervals for such an approach considerably under-estimate the intended coverage. We propose a novel and simple method to incorporate the uncertainty in the threshold selection in inference. The method is based on a double-bootstrap procedure, and we show that the coverage probabilities of confidence intervals from our approach are close to the required nominal levels, thus ensuring our inferences provide meaningful information for design policies.
@@
In Section 2 we illustrate problems with threshold selection and outline existing strategies. Section 3 describes the core existing automated methods while Section 4 introduces our procedure for the selection of a threshold, contrasting it with that of Varty et al., (2021). Section 5 then presents our proposed method for incorporating threshold uncertainty into return level inference. In Section 6, the proposed methods are compared against existing methods on simulated data. In Section 7, we apply our methodology to the widely studied troublesome dataset of the River Nidd, first analysed by Davison and Smith, (1990).
@@ 2 Background
For the Case 4 data, the plot shows that candidate thresholds above (below) $0.3$ are possibly appropriate (not appropriate) as CIs for higher candidate thresholds include (exclude) these estimates, and above 0.8 the point estimates appear more stable. Here $(u,\xi)=(1,0.1)$, so we can see that candidates below 0.3 are not suitable as $\xi$ is outside their CIs, but the true threshold is higher than may be selected using this plot. For the River Nidd, lower candidate threshold values imply a very heavy-tailed distribution ($\hat{\xi}\approx 0.5$), whilst high candidate thresholds imply a very short tail, with estimates ($\hat{\xi}\approx-1$) indicating the distribution’s upper end point is essentially the largest observation. As a result of this unusual behaviour, the Nidd data has become the primary example for non-trivial threshold selection (Davison and Smith,, 1990; Northrop and Coleman,, 2014). We apply our new method to this dataset in Section 7. Further examples are given in supplementary material S:2.2.
@@
Scarrott and MacDonald, (2012) and Belzile et al., (2023) review the extensive literature that aim to improve upon parameter stability plots. The latter characterises these methods, with a core reference, as follows: penultimate models (Northrop and Coleman,, 2014), goodness-of fit diagnostics (Bader et al.,, 2018), sequential-changepoint approaches (Wadsworth,, 2016), predictive performance (Northrop et al.,, 2017), and mixture models (Naveau et al.,, 2016). It also discusses semi-parametric inferences (Danielsson et al.,, 2001), but it excludes the development by Danielsson et al., (2019), with similarities to the goodness-of-fit approaches.
@@ 3 Existing automated threshold selection methods
Automated threshold selection methods choose from the set of candidate thresholds based on the evaluation of some metric. We comment on the two existing approaches, which we find to perform best, namely Wadsworth, (2016) and Northrop et al., (2017), with further details given in supplementary material S:3 and illustrations given in Section 6,
Wadsworth, (2016) addresses the dependence between MLEs of $\xi$, denoted by $\hat{\bm{\xi}}$, over candidate thresholds. Using asymptotic theory for the joint distribution of MLEs from overlapping samples of data, $\hat{\bm{\xi}}$ are transformed to the vector $\hat{\bm{\xi}}^{*}$ of normalised increments between successive $\hat{\bm{\xi}}$ values. For GPD data, asymptotically $\hat{\bm{\xi}}^{*}$ would be iid realisations from a univariate standard normal distribution, whereas if the data above any candidate threshold were not from a GPD, then the associated elements of $\hat{\bm{\xi}}^{*}$ would be better approximated by a non-standard normal distribution. This changepoint behaviour is used to select the threshold. The use of asymptotics leads to the choice of candidate thresholds causing either considerable sensitivity or failure of the method to converge, with both exacerbated by small samples. We also identify problems with the open source software, with failures when $\xi<0$.
Northrop et al., (2017) model data using the binomial-GPD (BGPD) model, which is GPD above $u$, with $\lambda_{u}=\mathbb{P}(X>u)$ a model parameter, and an improper uniform density, of value $1-\lambda_{u}$, below $u$. They use a Bayesian inference, for each candidate threshold, and assess the predictive density of GPD fits above a fixed validation threshold $v$, where $v$ is the largest candidate threshold. The selected threshold maximises the predictive ability of this model, above $v$, using leave-one-out cross-validation. The method is sensitive to the validation and candidate threshold set and to the prior joint density of the BGPD parameters.
@@ 4.1 Metric Choice
We propose the Varty et al., (2021) approach, and an adaptation of it, to select a threshold $u$, above which a GPD is consistent with a sample of size $n$ of iid, non-missing data from a continuous random variable. Both methods use a QQ-plot-based metric, but compare fit on different scales. Our method, which uses data on the original marginal scale, will be referred to as the expected quantile discrepancy (EQD) while the method of Varty et al., (2021), which transforms data to an Exponential$(1)$ marginal scale, will be termed the Varty method.
@@
The following makes the difference between the two methods precise, with both based on two tuning parameters $(k,m)$ which are defined below with default values discussed in Section 4.2. Let $\bm{x}_{u}=(x_{1},\ldots,x_{n_{u}})$ be the sample of $n_{u}$ excesses of candidate threshold $u$. The sample quantile function $Q(\cdot,\bm{x}_{u}):[0,1]\rightarrow\mathbb{R}^{+}$ is defined as the linear interpolations of the points $\left\{\left((i-1)/(n_{u}-1),{x}^{(i)}_{u}\right):i=1,\ldots,n_{u}\right\}$, where $x^{(i)}_{u}$ is the $i^{\text{th}}$ order statistic of $\bm{x}_{u}$ (increasing with $i$). The transformation to Exponential(1) margins is defined by $T(x;\sigma,\xi)=F^{-1}_{\text{Exp}}\{H(x;\sigma,\xi)\}$ where $F^{-1}_{\text{Exp}}$ is the inverse distribution function of an Exponential(1) variable and let $T(\bm{x}_{u};\sigma_{u},\xi)=\{T(x_{1};\sigma_{u},\xi),\ldots,T(x_{n_{u}};\sigma_{u},\xi)\}$. To incorporate the effect of sampling variability in the data into the threshold choice, the expected (average) deviation over the QQ-plot, calculated for the probabilities $\{p_{j}=j/(m+1):j=1,\dots,m\}$, is calculated across bootstrapped samples of $\bm{x}_{u}$, denoted $\bm{x}^{b}_{u}$ for the $b^{\text{th}}$ bootstrap sample, $b=1,\ldots,k$. For both methods, this results in the overall measure of fit $\hat{d}_{E}(u)=\sum_{b=1}^{k}d_{b}(u)/k$, where
$d_{b}(u)=\begin{cases}\frac{1}{m}\sum\limits_{j=1}^{m}\mid\frac{\hat{\sigma}^{b}_{u}}{\hat{\xi}^{b}_{u}}[(1-p_{j})^{-\hat{\xi}^{b}_{u}}-1]-Q(p_{j},\bm{x}^{b}_{u})\mid&\text{EQD}\\ \frac{1}{m}\sum\limits_{j=1}^{m}\mid-\log(1-p_{j})-Q(p_{j},\hat{T}(\bm{x}^{b}_{u};\hat{\sigma}^{b}_{u},\hat{\xi}^{b}_{u}))\mid&\text{Varty},\end{cases}$ (2)
@@
where $(\hat{\sigma}^{b}_{u},\hat{\xi}^{b}_{u})$ are the estimated GPD parameters fitted to the bootstrapped sample $\bm{x}^{b}_{u}$. The selected threshold minimises $\hat{d}_{E}$ over the set of candidate thresholds.
Now, we explain why the $\hat{d}_{E}(u)$ metric, for either the EQD or the Varty method, is helpful for assessing the validity of a candidate threshold $u$. For a given $u$, $d_{b}(u)$ evaluates the mean-absolute deviation between the $b^{\text{th}}$ bootstrap sample quantiles and the fitted model-based quantiles, i.e., the mean-absolute deviation from the line of equality in a QQ-plot for that particular bootstrap sample. This type of assessment by itself is not radical, as for any observed sample data, QQ-plots are the standard method of assessing model fit (Coles,, 2001) and Danielsson et al., (2019) minimise the maximum deviation of this plot. However, the key change in approach here, which has not been considered in any of the existing methods, is to take the average of such a metric over bootstrapped samples, with the less important novel feature being to interpolate the QQ-plot, using standard quantile definitions. The first change accounts for the sampling variability in the observed data, thus avoiding over-fitting of the GPD model to the observed data set which leads to higher threshold choices than necessary, reduced number of exceedances, and extra uncertainty in parameter and quantile estimates. The second change improves our approximation of the QQ error evaluation.
There are many ways to construct the metric in expression (2). We focus on the mean-absolute deviation on the QQ scale as Varty et al., (2021) found that this was more effective than using the mean-squared deviation on that scale and either metric in the PP scale. Although, higher (lower) sample quantiles exhibit greater (less) sampling variability, we choose not to weight values across the quantiles, as effectively we achieve an appropriate weighting by taking the $\{p_{j}\}$ values to be equally-spaced, as this leads to dense evaluation for lower sample quantiles and more sparse evaluation in the upper tail. Finally, we explore using the observed excesses within the sample quantile calculation, i.e., using $Q(p_{j},\bm{x}_{u})$ rather than $Q(p_{j},\bm{x}^{b}_{u})$ in metric (2), but results provided in supplementary material S:5.3 show benefit to this. Moreover, comparing observed sample quantiles against model-based quantiles obtained from a fit to a different (bootstrapped) sample would be unconventional.

### v4

@@ 1 Introduction
An inherent challenge in risk modelling is the estimation of high quantiles, known as return levels, beyond observed values. Such inference is important for designing policies or protections against future extreme events, e.g., in finance or hydrology (Smith,, 2003; Coles et al.,, 2003). Extreme value methods achieve this extrapolation by using asymptotically exact models to approximate the tail of a distribution above a high, within-sample, threshold. They require the selection of a threshold $u$, such that above $u$, extreme value methods provide a suitable model. Here, we develop novel methods for automatic selection of the threshold and for propagating the uncertainty in this selection into return level inferences.
Throughout, we assume that all data are realisations of an independent and identically-distributed (iid) univariate continuous random variable $X$ with unknown distribution function $F$, with upper endpoint $x^{F}:=\text{sup}\{x:F(x)<1\}$. Under weak conditions, Pickands, (1975) shows that for $X>u$, with $u<x^{F}$, the distribution of the rescaled excess $Y=X-u$, converges to the generalised Pareto distribution (GPD) as $u\rightarrow x^{F}$. To use this limit result in practice, a within-sample threshold $u$ is chosen, above which this limit result is treated as exact. Specifically, whatever the form of $F$, the excesses $Y$ of $u$ are modelled by the single flexible GPD$(\sigma_{u},\xi)$ family, with distribution function
@@
with $y>0$, $w_{+}=\text{max}(w,0)$, $(\sigma_{u},\xi)\in\mathbb{R}_{+}\times\mathbb{R}$ being scale and shape parameters. The exponential distribution arises when $\xi=0$, i.e., as $\xi\rightarrow 0$ in distribution (1), whereas for $\xi>0$, the distribution tail decay is polynomial. For $\xi<0$, $X$ has a finite upper end-point at $u-\sigma_{u}/\xi$ but is unbounded above for $\xi\geq 0$. To estimate the $(1-p)$th quantile, $x_{p}$, of $X$, for $p<\lambda_{u}:=\mathbb{P}(X>u)$, we can solve $\hat{F}(x_{p})=1-p$, where $\hat{F}(x_{p})=1-\hat{\lambda}_{u}[1-H(x_{p}-u;\hat{\sigma}_{u},\hat{\xi})]$, $\hat{\lambda}_{u}$ is the proportion of the realisations of $X$ exceeding $u$ and $(\hat{\sigma}_{u},\hat{\xi})$ are maximum likelihood estimates (MLEs) obtained by using realisations of the threshold excesses. Davison and Smith, (1990) overview the properties of the GPD.
@@
Threshold selection involves a bias-variance trade-off: too low a threshold is likely to violate the asymptotic basis of the GPD, leading to bias, whilst too high a threshold results in very few threshold excesses with which to fit the model, leading to large parameter and return level uncertainty. Thus, we must choose as low a threshold as possible subject to the GPD providing a reasonable fit to the data. There are a wide variety of methods aiming to tackle this problem (Scarrott and MacDonald,, 2012; Belzile et al.,, 2023) with the most commonly used methods suffering from subjectivity and sensitivity to tuning parameters.
@@
A novel automated approach to threshold selection is introduced by Varty et al., (2021) specifically for modelling large, human-induced earthquakes. These data are complex due to improvements in measurement equipment over time. The major implication of such change is that data are missing-not-at-random, with the dataset appearing to be realisations of a non-identically distributed variable, requiring a threshold $u(t)$ which varies with time $t$, even though the underlying process is believed to be identically distributed over $t$. Since excesses of $u(t)$ do not have the same GPD parameters over time, Varty et al., (2021) transform these to a common standard exponential distribution via the probability integral transform, using estimates of $(u(t),\sigma_{u(t)},\xi)$. They then quantify the model fit using a metric based on a QQ-plot and select a time-varying threshold that optimizes this metric. The key novel aspect of their assessment is the use of bootstrapping methods in the metric evaluation which fully accounts for the uncertainty in the GPD fit, which varies across threshold choices.
@@
Due to the lack of existing threshold selection methods designed for the context of Varty et al., (2021), that paper focuses on the data analysis rather than investigating the performance of the threshold selection method. We explore how their ideas can be best adapted to threshold selection in a univariate, iid data context. We find that a variant of the Varty et al., (2021) metric improves the performance and leads to substantially better results than existing automated methods, including greater stability with respect to tuning parameters.
@@
We differ from Varty et al., (2021) as we study both threshold selection and return level estimation when the truth is known. We also address an entirely different problem of how to incorporate the uncertainty resulting from threshold selection into return level estimation. Existing methods typically treat the threshold, once it has been selected, as known, for subsequent return level inference. The available data above candidate threshold choices are often few and so inference can be highly sensitive to the chosen threshold. Reliance on a single threshold leads to poor calibration of estimation uncertainty and as a result, can mislead inference. In particular, we show that the resulting confidence intervals for such an approach considerably under-estimate the intended coverage. We propose a novel and simple method, based on a double-bootstrap procedure, that incorporates the uncertainty in the selected threshold during inference. We show that the coverage probabilities of confidence intervals from our approach are close to the required nominal levels, thus ensuring our inferences provide meaningful information for design policies.
Ultimately, our aim is to provide a threshold selection method that does not require any user decisions to achieve adequate results. For example, the method should not be sensitive to the choice of candidate threshold grid, it should not require the estimation of a mode to select this grid, it should not have a limit on the number of candidate thresholds for a given sample size, nor should it exclude the possibility that the available data have been pre-processed, such as containing only the exceedances of some arbitrary level.
@@
In Section 2, we illustrate problems with threshold selection and outline existing strategies. Section 3 describes the core existing automated methods while Section 4 introduces our procedure for the selection of a threshold, contrasting it with that of Varty et al., (2021). Section 5 presents our proposed method for incorporating threshold uncertainty into return level inference. In Section 6, the proposed methods are compared against existing methods on simulated data. In Section 7, we apply our methodology to the widely studied troublesome dataset of the River Nidd, first analysed by Davison and Smith, (1990).
@@ 2 Background
For the Case 4 data, the plot shows that candidate thresholds above (below) $0.3$ are possibly appropriate (not appropriate) as CIs for higher candidate thresholds include (exclude) the corresponding shape parameter estimates, and above 0.8 the point estimates appear more stable. Here $(u,\xi)=(1,0.1)$, so we can see that candidates below 0.3 are not suitable as $\xi$ is outside their CIs, but the true threshold is higher than may be selected using this plot. For the River Nidd, lower candidate threshold values imply a very heavy-tailed distribution ($\hat{\xi}\approx 0.5$), whilst high candidate thresholds imply a very short tail, with estimates ($\hat{\xi}\approx-1$) indicating that the distribution’s upper end point is essentially the largest observation. As a result of this unusual behaviour, the Nidd data has become the primary example for non-trivial threshold selection (Davison and Smith,, 1990; Northrop and Coleman,, 2014). We apply our new method to this dataset in Section 7. Further examples of the problems encountered when using parameter stability plots are given in supplementary material S:2.2.
@@
Scarrott and MacDonald, (2012) and Belzile et al., (2023) review the extensive literature that aims to improve upon parameter stability plots. The latter characterises these methods, with a core reference, as follows: penultimate models (Northrop and Coleman,, 2014), goodness-of fit diagnostics (Bader et al.,, 2018), sequential-changepoint approaches (Wadsworth,, 2016), predictive performance (Northrop et al.,, 2017), and mixture models (Naveau et al.,, 2016). It also discusses semi-parametric inferences (Danielsson et al.,, 2001), but it excludes the development by Danielsson et al., (2019), with similarities to the goodness-of-fit approaches.
@@ 3 Existing automated threshold selection methods
Automated threshold selection methods aim to remove subjectivity from the choice of threshold by selecting an optimal threshold from a set of user-defined candidate thresholds based on optimising some criterion. We outline and compare the approaches of Wadsworth, (2016) and Northrop et al., (2017), which we find to perform best of the considered existing methods. Further details of these methods are given in supplementary material S:3.
Wadsworth, (2016) addresses the dependence between MLEs of $\xi$, denoted by $\hat{\bm{\xi}}$, over candidate thresholds. Using asymptotic theory for the joint distribution of MLEs from overlapping samples of data, $\hat{\bm{\xi}}$ are transformed to the vector $\hat{\bm{\xi}}^{*}$ of normalised increments between successive $\hat{\bm{\xi}}$ values. For GPD data, asymptotically, $\hat{\bm{\xi}}^{*}$ would be iid realisations from a standard normal distribution, whereas if the data above any candidate threshold were not from a GPD, the associated elements of $\hat{\bm{\xi}}^{*}$ would be better approximated by a non-standard normal. This changepoint behaviour is used to select the threshold. The underlying asymptotic arguments can cause considerable threshold sensitivity and the failure of the method to converge. Both issues are exacerbated by small samples and we identify systematic failures of the associated open source software when $\xi<0$. To reduce such problems, Wadsworth, (2016, Table 1) provides guidance on the number of candidate thresholds for a given sample size.
Northrop et al., (2017) model data using the binomial-GPD (BGPD) model, which is GPD above $u$, with $\lambda_{u}=\mathbb{P}(X>u)$ a model parameter, and an improper uniform density, of value $1-\lambda_{u}$, below $u$. They use Bayesian inference and, for each candidate threshold, assess the predictive density of GPD fits above a fixed validation threshold $v$, where $v$ is the largest candidate threshold. The selected threshold maximises the predictive ability of this model, above $v$, using leave-one-out cross-validation. The method is sensitive to the validation and candidate threshold set and to the prior joint density of the BGPD parameters.
@@ 4.1 Metric choice
We propose an adaptation of the Varty et al., (2021) approach to identify the threshold $u$ for which the sample excesses, arising from iid and non-missing realisations of a continuous random variable, are most consistent with a GPD model. Both methods use a QQ-plot-based metric to approximate the integrated absolute error (IAE) between the quantiles of the model and the data-generating process. Our method, the expected quantile discrepancy (EQD), uses the data on the original scale. In contrast, the method of Varty et al., (2021) transforms the data to an Exponential$(1)$ marginal scale and will be termed the Varty method.
@@
The following makes the difference between the two methods precise. Let $\bm{x}_{u}=(x_{1},\ldots,x_{n_{u}})$ be the sample of $n_{u}$ excesses of candidate threshold $u$ and $\bm{q}=\{q_{i}=(i-1)/(n_{u}-1):i=1,\ldots,n_{u}\}$ be the vector of probability plotting points corresponding to the sample size of $\bm{x}_{u}$. The sample quantile function $Q(\cdot;\bm{x}_{u},\bm{q}):[0,1]\rightarrow\mathbb{R}^{+}$ is defined as the linear interpolations of the points $\left\{(q_{i},{x}^{(i)}_{u}):i=1,\ldots,n_{u}\right\}$, with $x^{(i)}_{u}$ denoting the $i^{\text{th}}$ order statistic of $\bm{x}_{u}$ (increasing with $i$), where any ties are handled similarly through linear interpolation. The transformation to Exponential(1) margins is defined by $T(x;\sigma,\xi)=F^{-1}_{\text{Exp}}\{H(x;\sigma,\xi)\}$ where $F^{-1}_{\text{Exp}}$ is the inverse distribution function of an Exponential(1) variable, $H$ is the GPD function (1), and let $T(\bm{x}_{u};\sigma_{u},\xi)=\{T(x_{1};\sigma_{u},\xi),\ldots,T(x_{n_{u}};\sigma_{u},\xi)\}$. To incorporate the effect of sampling variability in the data into the threshold choice, the expected (average) deviation over the QQ-plot, calculated for the probabilities $\{p_{j}=j/(m+1):j=1,\dots,m\}$, is calculated across bootstrapped samples of $\bm{x}_{u}$, denoted $\bm{x}^{b}_{u}$ for the $b^{\text{th}}$ bootstrap sample, $b=1,\ldots,B$. For both methods, this results in the overall measure of fit $\hat{d}_{E}(u)=\sum_{b=1}^{B}d_{b}(u)/B$, where
$d_{b}(u)=\begin{cases}\frac{1}{m}\sum\limits_{j=1}^{m}\mid\frac{\hat{\sigma}^{b}_{u}}{\hat{\xi}^{b}_{u}}[(1-p_{j})^{-\hat{\xi}^{b}_{u}}-1]-Q(p_{j};\bm{x}^{b}_{u},\bm{q})\mid&\text{EQD}\\ \frac{1}{m}\sum\limits_{j=1}^{m}\mid-\log(1-p_{j})-Q(p_{j};\hat{T}(\bm{x}^{b}_{u};\hat{\sigma}^{b}_{u},\hat{\xi}^{b}_{u}),\bm{q})\mid&\text{Varty},\end{cases}$ (2)
@@
and $(\hat{\sigma}^{b}_{u},\hat{\xi}^{b}_{u})$ are the estimated GPD parameters fitted to the bootstrapped sample $\bm{x}^{b}_{u}$. The selected threshold minimises the estimated IAE, $\hat{d}_{E}$, over a set of candidate thresholds. In Sections 4.2 and 4.3 respectively, we justify the choices made in the formulation of the EQD metric and discuss our recommendation for default values for the tuning parameters $(B,m)$.
In supplementary material S:5.2, we compare the EQD and Varty methods through an extensive simulation study to assess which version of metric (2) performs better for threshold selection and quantile estimation. For threshold selection, the methods perform similarly; each method narrowly achieves the smallest root-mean-square error (RMSE) in two of Cases 1-4, discussed in Section 6. However, for the estimation of high quantiles, the EQD outperforms the Varty method obtaining the lowest RMSE in the majority of cases and quantiles, due to having the smaller variance of estimates. We ultimately aim to estimate high quantiles accurately following threshold selection. Given that this study indicates that the EQD should be preferred for this aim and to avoid unnecessary repetition, we omit the results for the Varty method for the remainder of the paper.
4.2 Investigation of the EQD metric choice
For a given $u$, $d_{b}(u)$ evaluates the mean-absolute deviation between the $b^{\text{th}}$ bootstrap sample quantiles and the fitted model-based quantiles, i.e., the mean-absolute deviation from the line of equality in a QQ-plot for that particular bootstrap sample. This type of assessment by itself is not radical, as for any observed sample data, QQ-plots are the standard method of assessing model fit (Coles,, 2001). The novelty for assessing the validity of a candidate threshold $u$ comes from the way that the EQD metric is constructed.
There are a number of novel choices which we have made in the EQD metric that require justification, in particular; the use of the mean-absolute deviation; the choice of quantiles and their interpolation in the QQ-plot; the use of bootstrap samples; and that the observed data are not explicitly used in the metric. We examine each of these features in the supplementary material, through simulation studies involving the case studies of Section 6. For each feature, we find positive evidence for our selections. Below, we explain why we made these choices and outline how they performed relative to other alternative formulations.
We focus on the mean-absolute deviation on the QQ scale as Varty et al., (2021) found that this was more effective than using the mean-squared deviation on that scale and either metric on the PP scale. Our simulation studies also found this to be a more robust measure of fit than the maximum deviation, as proposed by Danielsson et al., (2019).
We choose to define $(p_{j},q_{i})$ as the standard plotting points for a QQ-plot assessment and to weight contributions to $d_{b}(u)$ equally across the quantiles. Although higher (lower) sample quantiles exhibit greater (less) sampling variability, equal weighting is appropriate when taking the $\{p_{j}\}$ values to be equally-spaced because for any $\xi>-1$, the GPD density is monotonically decreasing. This leads to dense evaluation for lower sample quantiles and more sparse evaluation in the upper tail. The choice of equal weighting on this scale is motivated and supported by empirical evidence in Varty et al., (2021). We considered other $(p_{j},q_{i})$ definitions but found that there was no systematic ordering of the performance over these definitions and any differences in RMSE for the thresholds selected by the EQD were minimal, especially when compared to the differences between the EQD and existing methods.
The average over bootstrapped samples in metric (2) is not a standard use of bootstrapping, i.e., we utilise the bootstraps in a measure of fit rather than to describe the uncertainty in some estimated quantity. Our aim in doing this is to account for the sampling variability in the observed data, thus avoiding over-fitting of the GPD model to the observed dataset which could lead to higher threshold choices than necessary, reduced numbers of exceedances, and extra uncertainty in parameter and quantile estimates. To confirm this, we considered using only the observed sample values in the metric. This leads to higher and more variable thresholds choices in a variety of cases and an overall performance which is either noticeably worse or at best, comparable to our approach.
One may also be concerned that $\bm{x}_{u}$ is not included directly in metric (2). To address this, we additionally explored the effect of using $Q(p_{j};\bm{x}_{u},\bm{q})$ instead of $Q(p_{j};\bm{x}^{b}_{u},\bm{q})$ within the EQD metric, despite it being unconventional to compare sample quantiles to those of a model fitted using a different (bootstrapped) sample. We found no benefit to doing so. Moreover, using only $\bm{x}_{u}$ to estimate the IAE ignores that this estimate would change for another realisation of the same data generating process and that variability in this estimate increases with $u$. Our approach utilises the bootstrap resamples in the measure of fit to provide more stability in the threshold choice and allow us to account for the increasingly uncertain parameter and quantile estimates as the threshold increases.